# Spectral-Operator Calculus (Part A): Trace-Form Evaluators and Spectral Growth Taxonomy

### v1

@@ 5. Examples
Taken together, these examples show how the general calculus specializes in concrete settings. Once the spectral geometry is fixed, every evaluator satisfying Axioms A1-A5 assigns values by integrating a single profile $h$ against the spectral measure, and the growth class of $D$ determines which transforms $f(D)$ fall into the trace-class envelope $\mathcal{C}_{1}$. In this sense the examples are not separate constructions but instances of the same mechanism: lawful scalar data on these models are fully dictated by spectral geometry and the fixed profile $h$, as encoded in the trace-form representation of Theorem~3.1.
@@ 5.1. Finite-rank diagonal operators
Under the hypotheses of Theorem~3.1, the evaluator $\mathcal{E}$ restricted to this finite-rank class must agree with a weighted spectral trace:
@@ 5.2. Diagonal models for growth classes
We next give basic diagonal examples realizing each of the growth classes introduced in Definition~3.4.
@@ 5.2.1. Polynomial growth
so $D_{\mathrm{poly}}\in\mathcal{G}_{\mathrm{poly}}$ with exponent $d=1$. Note that $D_{\mathrm{poly}}$ is not trace class, hence $D_{\mathrm{poly}}\notin\mathcal{C}_{1}$; it is used here purely as a growth model and not as an input for the trace-form representation of Theorem~3.1.
@@ 5.2.2. Stretched-exponential behavior
Let $\beta>0$ and define
$D_{\mathrm{str}}=\operatorname{diag}(n^{\beta})_{n\geq 1}.$
Then
$N_{D_{\mathrm{str}}}(\Lambda)=\#\{n\geq 1:n^{\beta}\leq\Lambda\}=\#\{n\geq 1:n\leq\Lambda^{1/\beta}\}\sim\Lambda^{1/\beta}.$
This is still polynomial growth; the example is convenient for testing the monotone functional calculus in Theorem~3.5.
@@ 5.2.3. Logarithmic or slower growth
as $\Lambda\to\infty$. Thus $D_{\log}$ has logarithmic spectral growth in the sense of Definition~3.4(3) and provides a basic model for the “logarithmic or slower” class $\mathcal{G}_{\log}$.
@@ 5.3. Evaluators on Laplacian-type spectra
Under the standing assumptions, Theorem~3.1 implies that any evaluator $\mathcal{E}$ on this class agrees with a weighted trace-form observable:
@@ 5.4. Stability under sums and tensor products
We illustrate the closure statements in Theorem~3.5 with simple constructions.
@@ 5.4.1. Orthogonal sums
Let $D_{\mathrm{poly}}$ be as in Section 5.2 and $E$ another diagonal operator with the same growth rate, for instance
@@ 5.4.2. Tensor products
for any $\varepsilon>0$ and some $C>0$, so $D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}$ lies in $\mathcal{G}_{\mathrm{poly}}$ with exponent at most $2$. This is consistent with the tensor-product estimate in Lemma~4.2.
@@ 5.4.3. Monotone transforms
so polynomial growth is preserved under such transforms. This is a concrete instance of the monotone calculus bound in Lemma~4.1.
These examples confirm, in explicit models, that the growth classes and the structural operations of the spectral calculus are compatible in the way asserted by Theorem~3.5.
@@ 5.5. Evaluator limits under dominated convergence
Assume in addition that each $f_{k}(D)$ and $f(D)$ lies in the trace-class envelope $\mathcal{C}_{1}$. Then Theorem~3.1 gives
@@
for the Borel-measurable nondecreasing transform $h$ and constant $c_{\mathcal{E}}>0$ provided by Theorem~3.1. Under the same domination hypotheses, dominated convergence on the spectral side yields
@@ 6.1. Consequences
Let $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\log}$ denote the polynomial, stretched-exponential, and logarithmic-or-slower growth classes defined in Definition 3.4. Then:
1. The families $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\log}$ are pairwise disjoint. 2. Each family is closed under monotone functional calculus by regularly varying transforms. 3. Each family is closed under finite orthogonal sums. 4. In the polynomial case, each family is closed under tensor products: if $D,E\in\mathcal{G}_{\mathrm{poly}}$, then $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
@@ 6.2. Limitations
  • Spectral representation.
    The calculus is formulated entirely in the setting of the spectral theorem: each model is encoded by a self-adjoint operator $D$ on a separable Hilbert space together with its spectral measure, and evaluators act on the bounded Borel transforms $f(D)$ obtained from the functional calculus. If a system is given only in terms of scalar data or some other non-operator description, it must first be realized in this self-adjoint spectral form before the evaluator framework applies; no claims are made for evaluators outside this operator-theoretic setting.
@@
    Theorem~3.1 (trace-form representation) is proved only for inputs in the trace-class envelope $\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$. Within $\mathcal{C}_{1}$, the axioms A1-A5 force any evaluator to be of trace form with a single Borel nondecreasing profile $h$ and an overall normalization $c_{\mathcal{E}}$; in particular, there is no additional scalar freedom once these choices are fixed. Outside $\mathcal{C}_{1}$ the calculus does not assert trace-form representation, and additional hypotheses would be needed to control evaluators on non-trace-class transforms.
@@
    Dependence on projector-locality.
    Theorem~3.1 (trace-form representation) relies critically on projector-locality. If $\mathcal{E}$ is only unitary invariant and extensive on orthogonal sums, but not projector-local, then there exist natural examples (e.g. operator norm on finite-rank self-adjoint operators) that cannot be written in trace form and violate the additivity required on spectral partitions. Such counterexamples are discussed in detail in Appendix C.
@@
    The spectral growth taxonomy in Definition~3.4 and Theorem~3.5 assumes that $D$ has discrete spectrum with finite multiplicities and that the counting function $N_{D}(\Lambda)$ satisfies the regularity conditions of Convention~2.2. For operators with continuous spectrum or highly degenerate spectral measures, a different notion of growth (e.g. via spectral density functions or integrated density of states) is required. These extensions are not treated here.
@@
    The closure of growth classes under monotone functional calculus uses regular variation for the transforming function $f$. If $f$ has oscillatory or highly irregular behavior at infinity, the simple asymptotic relations of Lemma~4.1 may fail, and the effect on growth class can be subtler. In such cases, additional hypotheses on $f$ are needed.
@@ 6.3. Falsifiable internal tests
Refutation strategy: Exhibit a specific $\mathcal{C}$, an evaluator $\mathcal{E}$, and a family of test pairs $(f,D)$ with $f(D)\in\mathcal{C}_{1}$ for which no choice of Borel nondecreasing $h$ and scalar $c_{\mathcal{E}}$ can satisfy the trace-form identity above. Any such example would contradict Theorem~3.1 and falsify the trace-form representation claim for that evaluator.
@@
Setup: Let $D$ lie in one of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\log}$. Consider monotone regularly varying transforms $f$, orthogonal sums $D\oplus E$, and tensor products $D\otimes E$ with partners $E$ in the same class.
Claim: The transformed operators remain in the same growth class and satisfy the asymptotic bounds stated in Theorem~3.5.
@@
Setup: Fix a self-adjoint $D$, a sequence of bounded continuous functions $f_{k}\to f$ uniformly, and a dominating spectral transform $Y$ as in Corollary~6.3.
@@ 7. Conclusion
The limitations recorded in Section 6 indicate where further hypotheses or different techniques would be required: relaxing projector-locality, dropping dominated continuity, treating continuous or highly degenerate spectra, or employing more aggressive regularizations lie beyond the scope of the present framework. Within its stated domain, however, the calculus developed here supplies a robust backbone for subsequent work. Later papers in this series take the trace-form evaluator and growth taxonomy as given and build on them to analyze admissible kernels, invariant trace cutoffs, analytic continuation of spectral flows, and boundary reconstruction problems in more elaborate operator-theoretic settings.
@@ Acknowledgements
for the trace-class envelope of $\mathcal{C}$, and restrict attention to positive trace-class inputs $X\in\mathcal{C}_{1}^{+}$. Extension to general self-adjoint trace-class operators follows by decomposing into positive and negative parts.
Our goal is to prove Theorem~3.1: on $\mathcal{C}_{1}^{+}$ the evaluator $\mathcal{E}$ is of trace form with a Borel-measurable nondecreasing transform $h$, unique up to a positive scalar factor.
@@ A.3. Normalization and uniqueness
We now match the representation of Lemma~A.4 with the statement of Theorem~3.1 and record uniqueness up to a scalar factor.
@@ 2. If $\tilde{h}$ is another Borel-measurable function and $\tilde{c}_{\mathcal{E}}>0$ satisfy
Setting $c_{\mathcal{E}}:=1$ in Lemma~A.4 yields the simplest normalization. The general form of Theorem~3.1, with arbitrary $c_{\mathcal{E}}>0$ and $h$ unique up to a positive scalar factor, is recovered by replacing $h$ with $ah$ and $c_{\mathcal{E}}$ with $1/a$.
@@ A.4. Measure-theoretic formulation (optional viewpoint)
For completeness, we record the scalar-measure formulation of the evaluator, which parallels the classical trace-form representation via Radon-Nikodym. This subsection is not required for the proof of Theorem~3.1 on $\mathcal{C}_{1}$, but clarifies the connection with spectral measures and addresses measure-theoretic concerns.
@@ Appendix B Counting-function asymptotics and growth stability
In this appendix we collect the counting-function and Tauberian estimates used in the spectral growth taxonomy of Definition~3.4 and Theorem~3.5. Throughout we work with self-adjoint operators with discrete spectrum and finite multiplicities, so that the eigenvalues can be listed in nondecreasing order and counted by a monotone step function.
@@ B.1. Counting functions, asymptotic notation, and regular variation
$0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots,\qquad\lambda_{n}\to\infty\text{ as }n\to\infty.$
@@
$N_{D}(\Lambda):=\#\{n\geq 1:\lambda_{n}\leq\Lambda\},\qquad\Lambda\geq 0.$ (B.1)
@@ B.2. Monotone functional calculus and counting functions
We now justify the counting-function relation used in Lemma~4.1 and the monotone functional-calculus part of Theorem~3.5.
Let $D$ and $N_{D}$ be as above, and let $f:[0,\infty)\to[0,\infty)$ be bounded on compact intervals and nondecreasing, with $f(\Lambda)\to\infty$ as $\Lambda\to\infty$. By the spectral theorem, $f(D)$ is self-adjoint with discrete spectrum
$\mu_{n}:=f(\lambda_{n}),\qquad n\geq 1,$
again listed in nondecreasing order with multiplicity. Define
$N_{f(D)}(\Lambda):=\#\{n\geq 1:\mu_{n}\leq\Lambda\},\qquad\Lambda\geq 0.$
@@
This lemma justifies the counting-function relation used in the monotone functional-calculus part of Theorem~3.5: under the discrete-spectrum hypotheses and the monotonicity of $f$, we may treat $N_{f(D)}(\Lambda)$ and $N_{D}(f^{-1}(\Lambda))$ as identical for asymptotic purposes.
When one prefers to state the result with an error term, Lemma~B.5 implies
@@
as $\Lambda\to\infty$, in the sense of Definition~B.2.

### v2

@@ 5. Examples
Taken together, these examples show how the general calculus specializes in concrete settings. Once the spectral geometry is fixed, every evaluator satisfying Axioms A1-A5 assigns values by integrating a single profile $h$ against the spectral measure, and the growth class of $D$ determines which transforms $f(D)$ fall into the trace-class envelope $\mathcal{C}_{1}$. In this sense the examples are not separate constructions but instances of the same mechanism: lawful scalar data on these models are fully dictated by spectral geometry and the fixed profile $h$, as encoded in the trace-form representation of Theorem 3.1.
@@ 5.1. Finite-rank diagonal operators
Under the hypotheses of Theorem 3.1, the evaluator $\mathcal{E}$ restricted to this finite-rank class must agree with a weighted spectral trace:
@@ 5.2. Diagonal models for growth classes
We next give basic diagonal examples realizing each of the growth classes introduced in Definition 3.4.
@@ 5.2.1. Polynomial growth
so $D_{\mathrm{poly}}\in\mathcal{G}_{\mathrm{poly}}$ with exponent $d=1$. Note that $D_{\mathrm{poly}}$ is not trace class, hence $D_{\mathrm{poly}}\notin\mathcal{C}_{1}$; it is used here purely as a growth model and not as an input for the trace-form representation of Theorem 3.1.
As a simple test of the monotone functional calculus in the polynomial regime, fix $\beta>0$ and set $f(t)=t^{\beta}$ on $0,\infty)$. Then
$f(D_{\mathrm{poly}})=\operatorname{diag}(n^{\beta})_{n\geq 1},$
and
$N_{f(D_{\mathrm{poly}})}(\Lambda)=\#\{n\geq 1:n^{\beta}\leq\Lambda\}=\#\{n\geq 1:n\leq\Lambda^{1/\beta}\}\sim\Lambda^{1/\beta}.$
In particular $f(D_{\mathrm{poly}})\in\mathcal{G}_{\mathrm{poly}}$, in agreement with Theorem 3.5(1).
@@ 5.2.2. Stretched-exponential behavior
Fix constants $c>0$ and $0<\alpha<1$, and define
$D_{\mathrm{str}}=\operatorname{diag}\Big(\Big(\tfrac{1}{c}\log(n+1)\Big)^{1/\alpha}\Big)_{n\geq 1}.$
Then for $\Lambda\geq 0$,
$N_{D_{\mathrm{str}}}(\Lambda)=\#\Big\{n\geq 1:\Big(\tfrac{1}{c}\log(n+1)\Big)^{1/\alpha}\leq\Lambda\Big\}=\#\{n\geq 1:n+1\leq\exp(c\,\Lambda^{\alpha})\}.$
Hence
$N_{D_{\mathrm{str}}}(\Lambda)\sim\exp(c\,\Lambda^{\alpha})\quad(\Lambda\to\infty),$
so $D_{\mathrm{str}}\in\mathcal{G}_{\mathrm{str}}$. Note that “stretched-exponential” refers to the asymptotic growth of the counting function $N_{D}$, not to the eigenvalue sequence $(\lambda_{n}(D))$ itself.
@@ 5.2.3. Logarithmic or slower growth
as $\Lambda\to\infty$. Thus $D_{\log}$ has logarithmic spectral growth in the sense of Definition 3.4(3) and provides a basic model for the “logarithmic or slower” class $\mathcal{G}_{\mathrm{log}}$.
@@ 5.3. Evaluators on Laplacian-type spectra
Under the standing assumptions, Theorem 3.1 implies that any evaluator $\mathcal{E}$ on this class agrees with a weighted trace-form observable:
@@ 5.4. Stability under sums and tensor products
We illustrate the closure statements in Theorem 3.5 with simple constructions.
@@ 5.4.1. Orthogonal sums
Let $D_{\mathrm{poly}}$ be as in Section 5.2, and let $E$ be another diagonal operator with the same growth rate, for instance
@@ 5.4.2. Tensor products
for any $\varepsilon>0$ and some $C>0$, so $D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}$ lies in $\mathcal{G}_{\mathrm{poly}}$ with exponent at most $2$. This is consistent with the tensor-product estimate in Lemma 4.2.
@@ 5.4.3. Monotone transforms
so polynomial growth is preserved under such transforms. This is a concrete instance of the monotone calculus bound in Lemma 4.1.
These examples confirm, in explicit models, that the growth classes and the structural operations of the spectral calculus are compatible in the way asserted by Theorem 3.5.
@@ 5.5. Evaluator limits under dominated convergence
Assume in addition that each $f_{k}(D)$ and $f(D)$ lies in the trace-class envelope $\mathcal{C}_{1}$. Then Theorem 3.1 gives
@@
for the Borel-measurable nondecreasing transform $h$ and constant $c_{\mathcal{E}}>0$ provided by Theorem 3.1. Under the same domination hypotheses, dominated convergence on the spectral side yields
@@ 6.1. Consequences
Let $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ denote the polynomial, stretched-exponential, and logarithmic-or-slower growth classes defined in Definition 3.4. Then:
1. The families $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ are pairwise disjoint. 2. Each family is closed under monotone functional calculus by regularly varying transforms. 3. Each family is closed under finite orthogonal sums. 4. In the polynomial case, each family is closed under tensor products: if $D,E\in\mathcal{G}_{\mathrm{poly}}$, then $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
@@ 6.2. Limitations
  • Spectral representation (SOC-I).
    SOC-I is formulated in the setting of the spectral theorem: each model is a self-adjoint operator D on a separable Hilbert space together with its projection-valued spectral measure, and evaluators act on bounded Borel transforms f(D). Other operator classes require different functional-calculus inputs and, crucially, a different realization of the localization (coarse-graining) gate; the sectorial/holomorphic model case is developed separately.
@@
    Theorem 3.1 (trace-form representation) is proved only for inputs in the trace-class envelope $\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$. Within $\mathcal{C}_{1}$, the axioms A1-A5 force any evaluator to be of trace form with a single Borel nondecreasing profile $h$ and an overall normalization $c_{\mathcal{E}}$; in particular, there is no additional scalar freedom once these choices are fixed. Outside $\mathcal{C}_{1}$ the calculus does not assert trace-form representation, and additional hypotheses would be needed to control evaluators on non-trace-class transforms.
@@
    Dependence on projector-locality (PVM-localized model case).
    Theorem 3.1 (trace-form representation) relies critically on projector-locality, i.e. additivity across spectral (PVM) partitions. If $\mathcal{E}$ is only unitary invariant and extensive on orthogonal sums, but not projector-local, then there exist natural examples (e.g. operator norm on finite-rank self-adjoint operators) that cannot be written in trace form and violate the additivity required on spectral partitions. Such counterexamples are discussed in detail in Appendix C. In the sectorial model case, this axiom is replaced by a scale-band locality principle rather than by PVM additivity.
@@
    The spectral growth taxonomy in Definition 3.4 and Theorem 3.5 assumes that $D$ has discrete spectrum with finite multiplicities and that the counting function $N_{D}(\Lambda)$ satisfies the regularity conditions of Convention 2.2. For operators with continuous spectrum or highly degenerate spectral measures, a different notion of growth (e.g. via spectral density functions or integrated density of states) is required. These extensions are not treated here.
@@
    The closure of growth classes under monotone functional calculus uses regular variation for the transforming function $f$. If $f$ has oscillatory or highly irregular behavior at infinity, the simple asymptotic relations of Lemma 4.1 may fail, and the effect on growth class can be subtler. In such cases, additional hypotheses on $f$ are needed.
@@ 6.3. Falsifiable internal tests
Refutation strategy: Exhibit a specific $\mathcal{C}$, an evaluator $\mathcal{E}$, and a family of test pairs $(f,D)$ with $f(D)\in\mathcal{C}_{1}$ for which no choice of Borel nondecreasing $h$ and scalar $c_{\mathcal{E}}$ can satisfy the trace-form identity above. Any such example would contradict Theorem 3.1 and falsify the trace-form representation claim for that evaluator.
@@
Setup: Let $D$ lie in one of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$. Consider monotone regularly varying transforms $f$, orthogonal sums $D\oplus E$, and tensor products $D\otimes E$ with partners $E$ in the same class.
Claim: The transformed operators remain in the same growth class and satisfy the asymptotic bounds stated in Theorem 3.5.
@@
Setup: Fix a self-adjoint $D$, a sequence of bounded continuous functions $f_{k}\to f$ uniformly, and a dominating spectral transform $Y$ as in Corollary 6.3.
@@ 7. Conclusion
The limitations recorded in Section 6.2 indicate where further hypotheses or different techniques would be required: relaxing projector-locality, dropping dominated continuity, treating continuous or highly degenerate spectra, or employing more aggressive regularizations lie beyond the scope of the present framework. Within its stated domain, however, the calculus developed here supplies a robust backbone for subsequent work. Later papers in this series take the trace-form evaluator and growth taxonomy as given and build on them to analyze admissible kernels and cutoff schemes, analytic continuation procedures, and boundary and reconstruction problems in more elaborate operator-theoretic settings, including sectorial/holomorphic models where localization is implemented on log-scale rather than by spectral projections.
@@ Acknowledgements
for the trace-class envelope of $\mathcal{C}$, and
$\mathcal{C}_{1}^{+}:=\{X\in\mathcal{C}_{1}:X\geq 0\}$
for its positive cone. We restrict attention to inputs $X\in\mathcal{C}_{1}^{+}$. Extension to general self-adjoint trace-class operators follows by decomposing into positive and negative parts.
Our goal is to prove Theorem 3.1: on $\mathcal{C}_{1}^{+}$ the evaluator $\mathcal{E}$ is of trace form with a Borel-measurable nondecreasing transform $h$, unique up to a positive scalar factor.
@@ A.3. Normalization and uniqueness
We now match the representation of Lemma A.4 with the statement of Theorem 3.1 and record uniqueness up to a scalar factor.
@@ 2. If $\tilde{h}$ is another Borel-measurable function and $\tilde{c}_{\mathcal{E}}>0$ satisfy
Setting $c_{\mathcal{E}}:=1$ in Lemma A.4 yields the simplest normalization. The general form of Theorem 3.1, with arbitrary $c_{\mathcal{E}}>0$ and $h$ unique up to a positive scalar factor, is recovered by replacing $h$ with $ah$ and $c_{\mathcal{E}}$ with $1/a$.
@@ A.4. Measure-theoretic formulation (optional viewpoint)
For completeness, we record the scalar-measure formulation of the evaluator, which parallels the classical trace-form representation via Radon-Nikodym. This subsection is not required for the proof of Theorem 3.1 on $\mathcal{C}_{1}$, but clarifies the connection with spectral measures and addresses measure-theoretic concerns.
@@ Appendix B Counting-function asymptotics and growth stability
In this appendix we collect the counting-function and Tauberian estimates used in the spectral growth taxonomy of Definition 3.4 and Theorem 3.5. Throughout we work with self-adjoint operators with discrete spectrum and finite multiplicities, so that the eigenvalues can be listed in nondecreasing order and counted by a monotone step function.
@@ B.1. Counting functions, asymptotic notation, and regular variation
$0\leq\lambda_{1}(D)\leq\lambda_{2}(D)\leq\cdots,\qquad\lambda_{n}(D)\to\infty\text{ as }n\to\infty.$
For brevity in this appendix we abbreviate $\lambda_{n}:=\lambda_{n}(D)$. (Since the spectrum is contained in $0,\infty)$, absolute values in counting conditions are redundant in this appendix.)
@@
$N_{D}(\Lambda):=\#\{n\geq 1:\lambda_{n}(D)\leq\Lambda\},\qquad\Lambda\geq 0.$ (B.1)
@@ B.2. Monotone functional calculus and counting functions
We now justify the counting-function relation used in Lemma 4.1 and the monotone functional-calculus part of Theorem 3.5.
Let $D$ and $N_{D}$ be as above, and let $f:[0,\infty)\to[0,\infty)$ be bounded on compact intervals and nondecreasing, with $f(\Lambda)\to\infty$ as $\Lambda\to\infty$. By the spectral theorem, $f(D)$ is self-adjoint with discrete spectrum $\{f(\lambda_{n})\}_{n\geq 1}$, again listed in nondecreasing order with multiplicity. Define
$N_{f(D)}(\Lambda):=\#\{n\geq 1:f(\lambda_{n})\leq\Lambda\},\qquad\Lambda\geq 0.$
@@
This lemma justifies the counting-function relation used in the monotone functional-calculus part of Theorem 3.5: under the discrete-spectrum hypotheses and the monotonicity of $f$, we may treat $N_{f(D)}(\Lambda)$ and $N_{D}(f^{-1}(\Lambda))$ as identical for asymptotic purposes.
When one prefers to state the result with an error term, Lemma B.5 implies
@@
as $\Lambda\to\infty$, in the sense of Definition B.2.